OpenStudy (anonymous):

Change the equation: 3x-y=3 to slope intercept form.

OpenStudy (anonymous):

What does slope intercept look like?

OpenStudy (anonymous):

y=mx + b

OpenStudy (anonymous):

Ok, so can you (using algebra) move the y to one side and everything else to the other?

OpenStudy (anonymous):

well, I got y=-3x-3 but I'm not sure if that's right.

OpenStudy (anonymous):

Nearly.

OpenStudy (anonymous):

\[3x-y = 3\] \[3x -y +y = 3 + y\] \[3x= 3 + y\] \[3x-3 = 3 -3+y\] \[3x-3 = y\] \[y=3x-3\]